A triangle is placed in a semicircle with a radius of 9 cm, as shown below. Find the area of the shaded region. Use the value for 3.14, and do not round your answer. Be sure to include the correct unit in your answer.
1
Expert's answer
2019-03-25T11:38:03-0400

S=Ssemicircle-Sstriagle=(pr^2)/2-(2r/2*r)=3.14*81/2-81=46.17cm^2